OpenStudy (anonymous):

Five litres of water wee poured from tank A to tank B and 10 litres were then poured from tank A to Tank C. If tank A originally had 10 more litres of water than tank C, how many more litres of water does tank C now have more than tank A?

OpenStudy (apoorvk):

Let the original no. of litres of water in tank A be 'x'. Then C originally had = x-10 Now, C has = 'x'. But A lost '10' to C and '5' to B as well, so A now has =x-10-5 = x-15 helps?

OpenStudy (apoorvk):

So, now A has 'x-15' and C has 'x'. That means C has '15' more than A. right?

OpenStudy (dumbcow):

A = 10+C A' = A-15 --> A = 15+A' C' = C+10 --> C = C'-10 by substitution 15+A' = 10+(C'-10) 15+A' = C' tank C now has 15 more than A